HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SI TRACEABLE SATURATED ABSORPTION SPECTROSCOPY OF OZONE AT THE 50 KHZ LEVEL USING A QUANTUM CASCADE LASER

AT 9.5 MICROMETRE

N. CAHUZAC, LERMA, Sorbonne Université, Observatoire de Paris, PSL Université, CNRS, FRANCE; LPL, Université Sorbonne Paris Nord, CNRS, FRANCE; H. ELANDALOUSSI, M. LEULIET, P. MARIE-JEANNE, C. ROUILLÉ, T. ZANON-WILLETTE, \underline{C. JANSSEN}, LERMA, Sorbonne Université, Observatoire de Paris, PSL Université, CNRS, FRANCE; M.

MANCEAU, O. LOPEZ, B. DARQUIÉ, LPL, Université Sorbonne Paris Nord, CNRS, FRANCE; M. ABGRALL, LNE-SYRTE, Sorbonne Université, Observatoire

de Paris, PSL Université, CNRS, FRANCE

Significant progress on the quantitative spectroscopy of ozone in the UV^{1}^{,}^{2}^{,}^{3} and the IR^{4}^{,}^{5}^{,}^{6}^{,}^{7}^{,}^{8} has been made recently, addressing the long-standing issue of the missing consistency between UV and IR spectroscopic data for the remote sensing of ozone.^{9} However, some uncertainty remains whether these improvements suffice to account for the striking observation of opposite trends in the global tropospheric ozone burden derived from satellite instruments (e.g. OMI or TROPOMI operating in

1doi:10.1088/1681-7575/ab0bdd, J. T. Hodges, J. Viallon, et al. Recommendation of a consensus value of the ozone absorption cross-section at 253.65 nm based on a literature review, Metrologia 56, 034001 (2019).

2doi:10.5194/amt-11-1707-2018, C. Janssen, H. Elandaloussi, J. Gröbner, A new photometric ozone reference in the Huggins bands: the absolute ozone absorption cross section at the 325 nm HeCd laser wavelength, Atmos. Meas. Tech. 11, 1707–1723 (2018).

3doi:10.5194/amt-13-5845-2020, J. Bak, X. Liu, M. Birk, G. Wagner, I. E. Gordon, K. Chance, Impact of using a new ultraviolet ozone absorption cross-section dataset on OMI ozone profile retrievals, Atmos. Meas. Tech. 13, 5845–5854 (2020).

4doi:10.1016/j.jqsrt.2021.107949, I. Gordon, L. Rothman, et al. The HITRAN2020 molecular spectroscopic database, J. Quant. Spectrosc. Radiat. Transfer 277, 107949 (2022).

5doi:10.1016/j.jqsrt.2021.108051, C. Janssen, C. Boursier, et al. Multi-spectral investigation of ozone: Part I. Setup & uncertainty budget, J. Quant. Spectrosc. Radiat. Transfer 279, 108051 (2022).

6doi:10.1016/j.jqsrt.2021.108050, D. Jacquemart, C. Boursier, H. Elandaloussi, P. Jeseck, Y. Té, C. Janssen., Multi-spectral investigation of ozone: Part II. Line intensity measurements at one percent accuracy around 5 μm and 10 μm, J. Quant. Spectrosc. Radiat. Transfer 279, 108050 (2022).

7doi:10.1063/1.5089134, V. G. Tyuterev, A. Barbe, D. Jacquemart, C. Janssen, S. N. Mikhailenko, E. N. Starikova, Ab initio predictions and laboratory validation for consistent ozone intensities in the MW, 10 and 5 μm ranges, J. Chem. Phys. 150 (2019).

8doi:10.1016/j.jqsrt.2021.107801, V. Tyuterev, A. Barbe, S. Mikhailenko, E. Starikova, Y. Babikov, Towards the intensity consistency of the ozone bands in the infrared range: Ab initio corrections to the S&MPO database, J. Quant. Spectrosc. Radiat. Transfer 272, 107801 (2021).

9doi:10.1021/jp0405411, B. Picquet-Varrault, J. Orphal, J.-F. Doussin, P. Carlier, J.-M. Flaud, Intercomparison of the ozone absorption coefficients in the mid-infrared (10 μm) and ultraviolet (300-350 nm) spectral regions, J. Phys. Chem. A 109, 1008 – 1014 (2005).

p-number: p131  Submitted on Wed, 14 Jun 2023 12:44:46 +0200

the UV versus IASI in the IR)^{10} or for the inconsistencies in the ozone retrieval from different IR bands using the same instrument.^{11} Unresolved issues in the IR spectral region at 9.5 μm are uncertainties and inconsistencies in current ozone broadening^{12} and pressure shift^{13} parameters, as well as the question of appropriately representing molecular line shapes.^{14} These issues are best investigated using SI traceable instrumentation which provides spectral resolution surpassing that of typical high resolution Fourier transform spectrometers.

Based on previous work at LPL,^{15}^{,}^{16} we have developed a new SI traceable high resolution laser spectrometer, which will be able to address these questions. A quantum cascade laser at 9.5 μm is frequency stabilized on an optical frequency comb itself stabilized to a 1.55 μm ultra-stable laser transmitted by fiber link through the REFIMEVE^{17} network from LNE-SYRTE. There it is measured against primary standards, thus assuring SI traceability of the frequency measurements. As a first demonstration of the spectrometer perfomances, we report on line positions in the Rbranch of the ozone ν_{3} band from saturated absorption spectroscopy with line center determinations at the 50 kHz level. This is almost three orders of magnitude more accurate than the current uncertainty index (< 10^{−3} cm^{−1}) given in the HITRAN 2020 database.

10doi:10.1525/elementa.291, A. Gaudel, O. R. Cooper, et al. Tropospheric ozone assessment report: Present-day distribution and trends of tropospheric ozone relevant to climate and global atmospheric chemistry model evaluation, Elementa: Sci. Anthrop. 6 (2018).

11doi:10.5194/amt-16-1443-2023, M. Kiefer, T. von Clarmann, et al. Version 8 IMK–IAA MIPAS ozone profiles: nominal observation mode, Atmos. Meas. Tech. 16, 1443–1460 (2023).

12doi:10.5194/amt-11-4707-2018, N. Glatthor, T. von Clarmann, G. P. Stiller, M. Kiefer, A. Laeng, B. M. Dinelli, G. Wetzel, J. Orphal, Differences in ozone retrieval in MIPAS channels A and AB: a spectroscopic issue, Atmos. Meas. Tech. 11, 4707–4723 (2018).

13doi:10.1016/j.jms.2017.12.009M. Minissale, T. Zanon-Willette, P. Jeseck, C. Boursier, C. Janssen. First pressure shift measurement of ozone molecular lines at 9.54 μm using a tunable quantum cascade laser. J. Molec. Spectrosc. 348, 103–113 (2018).

14doi:10.1016/j.jqsrt.2010.04.002, H. Tran, F. Rohart, C. Boone, M. Eremenko, F. Hase, P. Bernath, J.-M. Hartmann, Non-Voigt line-shape effects on retrievals of atmospheric ozone: Collisionally isolated lines, J. Quant. Spectrosc. Radiat. Transfer 111, 2012–2020 (2010).

15doi:10.1364/optica.6.000411, R. Santagata, D. B. A. Tran, et al. High-precision methanol spectroscopy with a widely tunable SI-traceable frequency-comb-based mid-infrared QCL, Optica 6, 411– 423 (2019).

16doi:10.1038/nphoton.2015.93, B. Argence, B. Chanteau, et al. Quantum cascade laser frequency stabilization at the sub-Hz level., Nature Phot. 9, 456–460 (2015).

17www.revimeve.fr
p-number: p131  Submitted on Wed, 14 Jun 2023 12:44:46 +0200